3. y - бу +9у 3 te3t У(0)%31,у (0)-4
Answer #1

y-Cy uqy : test Yna y-sytay=0 [y=ext -> Men - Gule dage - y= rett g= 8ext] - entf-68 +9)=0 ?.68+9=0 (8-3) 20 – 7:3 for2a,-0. 5 a 0 6ão 31 340 - Y6 335 - y - ynt yp I gettca test+ftet! Sycobsoj 1 = CU y(o)= c.exq, los e + 2 cope IT=C1] = y(t)